Suppose that the economy has entered a recession. What monetary policy action will a central bank take to restore full-employment output?


An increase in the price level will affect the money market and bond market in what ways?


If the interest rate on a one-year loan is 5% and the expected inflation rate is −2% for the same period, what is the expected real interest rate on the loan?


Spencer took a 9 percent one-year fixed-rate loan to buy a new car. He expected to pay a real interest rate of 5 percent. If at the end of the year Spencer only paid a 3 percent real interest rate, what is the actual or nominal interest rate?




1
Expert's answer
2019-11-26T11:23:54-0500

1.  If an economy is experiencing  recession threatens, the central bank uses an expansionary monetary policy to increase the supply of money, increase the quantity of loans, reduce interest rates, and shift aggregate demand to the right.

2.An increase in price level will cause decrease in money market. If there is a decrease in the money supply, bond prices will decrease. When the money supply decreases, nominal interest rates rises. Bond prices and interest rates are inversely related, so when interest rates go up, bond prices falls.

3.Real interest rate =nominal interest rate -inflation rate

                                 =5-(-2)=7%

4.nominal interest rate =real interest rate +inflation rate

                                      9 =3+x =6                                        

In this case, actual inflation rate is 6%